# The Peekaboo galaxy: new SALT spectroscopy and implications of archive HST data

### v1

@@ Abstract
Context. The dwarf galaxy Peekaboo (HIPASSJ1131–31) was recently identified as a Local Volume (LV) gas-rich and ’eXtremely Metal-Poor’ (XMP) dIrr. Its gas metallicity is ZZ⊙/50, with 1 uncertainty range of [Z⊙/72–Z⊙/35]). Its the ”Tip of Red Giant Stars” (TRGB)-based distance is of 6.80.7 Mpc. The Hubble Space Telescope (HST) data for its individual stars reveal that its older RGB stars comprise a smaller part, while the majority of visible stars have ages of less than one to a few Gyr. Thus, Peekaboo dwarf can be considered as the nearest record-low-Z dwarf. As such, the galaxy deserves a deeper multi-method study, including properties of both, young massive stars and the fainter older population, and its ionised gas and the dominant baryonic component of Hi gas.
Aims. We aim to obtain the higher S-to-N SALT optical spectra of Peekaboo two H ii regions, in order to improve the accuracy of its gas O/H and to determine abundances of Ne, S, N, Ar. With the archive HST images, we aim to identify the hot massive stars (including exciting the two H ii regions) and the XMP supergiants as important targets for the follow-up studies of their evolution with upcoming Extremely Large Telescopes.
Methods. We use the direct (Te) method for the east H ii region, in which the [Oiii]4363Å line is well detected, to estimate its parameter 12+log(O/H). Since in the west H ii region the line [Oiii]4363Å is not detected, its O/H is estimated via the empirical ”strong-line” method of Izotov et al. The resulting value of O/H is very close to that in the east H ii region.
Results. The new spectroscopy of Peekaboo dwarf allows us to improve substantially the accuracy of its direct O/H estimate, which appears of 12+log(O/H) = 6.990.06 dex. The new data reveal that emission lines in the E region consist of two components with the velocity difference of 65 km s-1. The fainter, approaching, component can be related to a fast-moving WR star thrown from a cluster or a binary system. Using the HST magnitudes and colour , we identify tentative O-type and very hot candidate WO stars, which are likely the ionising stars of the studied H ii regions.
Conclusions. With the new optical spectra, the Pekaboo galaxy is confirmed as the lowest-metallicity dwarf in the Local Volume and the valuable object for indeep multi-method studies. Its most luminous stars are separated for the follow-up ground-based brightness monitoring and spectroscopy.
@@
The ’local’ gas-rich XMP (eXtremely Metal-Poor) dwarfs are interesting by themselves since they challenge the modern cosmological models of galaxy formation and evolution, allowing to check their predictions for the formation and survival of objects with the extreme properties. The gas-rich XMP dwarfs on the sum of their observational properties look like galaxies in the early stages of their evolution.
@@
The existence of such unusual, very rare, so-called VYGs (Very Young Galaxies), was also predicted in N-body simulations (Tweed et al. 2018). VYGs are defined as galaxies formed the major part of their stellar mass within the last 1 Gyr. On the one hand, VYGs are expected to be very rare. On the other hand, their proportion in simulations is strongly depending on the type of Dark Matter (Cold DM versus Warm DM).
Besides the limitations on the cosmological models, which the statistics of XMP VYGs can imply, their study is of interest at least in two more areas.
@@
Firstly, due to the expected similarity of their properties to those of young galaxies in the early Universe (very gas-rich and very metal-poor). Being at the distances of the LV and its environs, they allow one to get a much deeper insights in their properties and to tie them with the processes of galaxy formation in the early Universe.
@@
The other important issue, related to stellar physics and galaxy evolution, is the basic understanding of formation and evolution of massive stars and their feedback into ISM in their parent galaxies (e.g. Eldridge, Stanway 2022). For cosmology of the early Universe, it is especially crucial to understand well properties and evolution of the massive stars with metallicities of tens - hundreds times lower than the solar one Z⊙. Such massive, very-low-Z stars can be found only in a handful XMP dwarf galaxies within the LV and its environs (Garcia et al. 2021). If identified, such XMP massive stars can be the primary targets for spectroscopic studies with the next generation ground-based telescopes. Meantime, the first steps in this direction are already presented for a nearby dIrr Leo A with Z 0.03 Z⊙ by Gull et al. (2022) and for four low-metallicity galaxies, including Leo P with Z 0.025 Z⊙ by Telford et al. (2024).
@@
The Pekaboo galaxy, the optical counterpart of Hi source HIPASS J1131–31, got its name since it was hidden within the halo of a nearby bright star (Koribalski et al. 2018). The galaxy was recently described and studied by Karachentsev et al. (2023). Referring this paper, ”Peekaboo dwarf is a faint (MV = –11.27 mag), the Local Volume (LV) gas-rich dIrr at the TRGB-based distance of 6.8 Mpc (distance module = 29.2 mag). Its gas metallicity, derived with the direct Te method, appeared one of the lowest of the known so far (12+log(O/H) = 6.990.16 dex)”. The cited uncertainties of O/H correspond to the range of a factor of two: [Z⊙/72–Z⊙/35]). The authors also discuss the probable ”youth” of this galaxy, paying attention to the abnormally small population of aged RGB stars.
@@
In this Letter, using the new SALT spectroscopy, we update the estimate of Peekaboo metallicity, increasing substantially its accuracy. Also, we use the archive HST images and the Extragalactic Distance Database (EDD)111https://edd.ifa.hawaii.edu photometry of its stars to identify the hot massive stars, ionising the two H ii regions, and several the brightest supergiants. The latter can be attractive targets for HST and ELT spectral study as the massive evolved stars with the record-low metallicities.
@@
The layout of the Letter is as follows. In Sect. 2, the data and analysis are briefly described. Sect. 3 presents the main results. They are discussed in Sect. 4 and summarised in Sect. 5. The details of data and results are shown in Appendices. The adopted scale is 34 pc in 1 arcsec.
@@ 2.1 SALT observations and analysis
We carried out long-slit optical spectroscopic observations of Peekaboo galaxy using the Robert Stobie Spectrograph (RSS; Burgh et al. 2003; Kobulnicky et al. 2003) installed at the Southern African Large Telescope (SALT; Buckley, Swart & Meiring 2006; O’Donoghue et al. 2006). Different VPH gratings were used with the long slit of 1.5″ by 8′ to cover the full spectral range from 3600 Å to 7250 Å. The position angle for all observations was 87.88° and slit position was exactly the same as it was shown in Figure 6 of Karachentsev et al. (2023), where reference star was used to put the slit in exactly the same position for different observations. All details are shown in Table 1, where all dates, seeing, covered spectral range, spectral scale and FWHM are shown. The final spectrum taken with grating PG0900 was used as a basic one, since it covers the most spectral range. Spectral data obtained using other gratings, were processed independently and emission lines were measured in the same spatial areas. After that, the line fluxes, measured on the spectra obtained with other gratings, were recalculated to the flux system obtained using the PG900 grating. Namely, the total fluxes of the bright lines, measured on spectra for both gratings, were used to determine the conversion factor, and then, the fluxes for the weaker lines were recalculated using this factor. Spectrophotometric standards were observed at twilight as part of the SALT standard calibration plan.
@@
Given that SALT is equipped with an Atmospheric Dispersion Compensator (ADC), the effect of atmospheric dispersion at an arbitrary long-slit is negligible. Additionally, the SALT design entails the movement of the telescope’s pupil during the track and exposures, thereby resulting in a constant alteration of the telescope’s effective area. Therefore, accurate absolute photometry and spectrophotometry are not possible. However, the relative flux calibration could be used. Thus, the relative energy distribution in the spectra could be obtained with SALT data.
@@
Fortunately, we have on the slit the reference star with the known magnitudes in several bands from the Legacy DR10 database Li et al. (2024), which we can use as a local standard. to derive the line fluxes and luminosity of H. See details in Appendix, Sect. B.
@@
The primary data reduction was done with the SALT science pipeline (Crawford et al. 2010). The long-slit reduction was done with the RSS pipeline described by Kniazev (2022). The 1D spectra extraction, emission lines measurement and calculation of chemical abundances was done in the way described by Kniazev et al. (2008). The fully reduced 2D spectra of Peekaboo are shown Figure 2 with labelling of the most prominent lines. Since the RSS detector is a mosaic and consists of three CCDs, the positions of two gaps between these CCDs are also shown.
@@
Errors in lines intensities have been propagated in the calculations of the reddening correction, electron temperatures and densities and have been propagated to the elemental abundance errors. The observed emission line intensities relatively to the F(H) line and the intensities corrected for interstellar extinction and underlying stellar absorption are presented in Table 2. The equivalent width (H) of the H line, the calculated absorption equivalent widths (abs) of the Balmer lines, the extinction coefficient (H) (a sum of the internal extinction in Peekaboo and foreground extinction in the Milky Way) and the subsequently recalculated value are also listed there.
@@ 2.2 Hubble Space Telescope data
HST data are well described by Karachentsev et al. (2023). Here we use the information on positions, magnitudes and colours of the hottest stars as the potential exciting stars of the observed H ii regions. The evolved luminous stars (supergiants) are also identified with this data. They should be the valuable material for a more advanced study of the lowest metallicity massive stars with the upcoming new facilities, including the ELT.
In Figure 1 we show two zoomed HST images. In the top panel, we show image in F606W filter, where all stars are marked with their numbers in Table 4. These is the same set of stars as in paper by Karachentsev et al. (2023). In the bottom panel we present the similarly zoomed colour HST image of Peekaboo, taken from the HST archive (Prog. SNAP 15992, PI R.B. Tully).
@@ 3.1 Structure and kinematics of the main H ii regions
Analysis of the long-slit spectra revealed the presence of two distinct H ii regions. Figure 3 shows the profile of the H line along the slit, based on the spectral data obtained with the PG900 grating. The analysis clearly shows two regions, which we will refer to as regions E and W, since the slit was oriented roughly in the east-west direction across the Galaxy. These regions are prominent as two peaks in the H profile, although they are less pronounced in the continuum profile. Region W is fainter, with an extent of about 3 arcsec, while the brighter region E extends over about 5 arcsec. We defined approximate boundaries for these regions (as shown in the figure) in order to extract 1d spectra along the slit. The aim was to minimise spectral contamination between the regions while maximising the useful signal from the emission lines. The selected region boundaries are also shown in the Figure 2 as red horizontal lines and in the top panel of Figure 1 as red circles centred on the reference star with the corresponding radii. These boundaries were applied consistently to all the acquired spectral data. The one-dimensional spectra obtained by summing within these limits are shown in Figures 4 and 5.
A more detailed analysis of the extracted one-dimensional spectra uncovers an interesting peculiarity. Namely, all observed emission lines in the spectrum of region E appear double-component. Every line exhibits an asymmetric blue wing. To test the hypothesis that all lines represent two velocity components, we first fit each line with a single Gaussian, and then – with a two-Gaussian model. The related improvement of the reduced statistic is evaluated. The improvement is very significant. For various lines, it varies by a factor of 2 to 6.5. Examples of the fitting for H and [O iii]Å lines and H and H, for gratings PG900 and PG3000 are presented in Figures 6 and 7, respectively.
We calculate the radial velocities of two components in the E region and that of the W region, taking the weighted means on the three independent measurements with grisms PG900, PG1800 and PG3000. Each of these velocities, in turn, was derived from the redshifts of 6–7 the strongest emission lines. The final values of Vhel for these three subsystems are as follows: East(red): 72110 km s-1; East(blue): 6563 km s-1, West: 6976 km s-1. Thus, we accept for the further discussion these three Peekaboo ionised gas systems, of which East(red) and East(blue) are co-spatial within the limits of our modest spatial resolution.
@@ 3.2 New estimates of O/H and other elements
The measured and corrected line fluxes and the derived element abundances are presented in Tables 2 and 3. Here we briefly summarise the obtained results. In Sect. 3.1 we found that there are three separate subsystems on the slit.
For the East(red) subsystem, the line [O iii]4363 is detected at the level of 6.5. This allows us to directly determine electron temperature Te and derive Oxygen and the other element abundances via the direct method. We obtain its 12+log(O/H)(Te) = 6.990.06 dex. The two non-direct methods, described in Sect. D, give for the same subsystem the O/H values within 0.02–0.03 dex of the former. For the West and the East(blue) subsystems, the line [O iii]4363 is not detected. Their O/H are derived with the ’strong-line’ method of Izotov et al. (2019). For both objects, the derived O/H is within 0.02–0.03 dex from (O/H)(Te) of the East(red).
3.3 HST data on XMP stars in Peekaboo dwarf
Having in the EDD (Anand et al. 2021) the table with parameters of all 56 individual stars within the Peekaboo dwarf body, we sort them on the colour and separate the hottest stars with –0.25 (see Sect. 4.2), expected to ionise the surrounding gas. We then look for such stars which fall within (or near) the borders of the E and W H ii regions. They are the most likely stars, exciting these H ii regions.
There are also several supergiants in this XMP dwarf, which also are very interesting, since they are the nearest massive evolved stars with the record-low metallicity. They are attractive targets for the spectral study with the HST and the next generation ELTs. See top panel of Fig. 1 with the HST image of Peekaboo with positions of these stars indicated.
The majority of the rest stars with of 0.0–0.5 mag and M –2.5 mag, most probably belong to the Blue Helium Burning (BHeB) type with ages less than 200 Myr (McQuinn et al. 2012).

### v2

@@ Abstract
Context. The dwarf galaxy Peekaboo (HIPASSJ1131–31) was recently identified as a local volume (LV) gas-rich and extremely metal-poor (XMP) dIrr. Its gas metallicity is ZZ⊙/50, with a 1 uncertainty range of [Z⊙/72–Z⊙/35]). Its ’tip of the red-giant branch’ distance is 6.80.7 Mpc. The Hubble Space Telescope (HST) data for its individual stars revealed that its older red-giant-branch stars comprise a smaller part of the galaxy, while the majority of visible stars have ages of less than one to a few gigayears. Thus, the Peekaboo dwarf can be considered as the nearest record-low Z dwarf. As such, the galaxy deserves a deeper multi-method study that examines the properties of its young massive stars and the fainter older population as well as its ionised gas and the dominant baryonic component of Hi gas.
Aims. We aim to obtain the higher S-to-N SALT optical spectra of two H ii regions in Peekaboo in order to improve the accuracy of its gas O/H and to determine abundances of Ne, S, N, and Ar. With archive HST images, we aim to identify the hot massive stars, including exciting the two H ii regions (i.e. east and west), and the XMP supergiants as important targets for follow-up studies of their evolution with upcoming extremely large telescopes.
Methods. We used the direct (Te) method for the east H ii region in which a [Oiii]4363Å line is well detected in order to estimate its parameter 12+log(O/H). In the west H ii region, the line [Oiii]4363Å is not detected, so we estimated its O/H via the empirical ’strong-line’ method of Izotov et al. The resulting value of O/H is very close to that in the east H ii region.
Results. The new spectroscopy of the Peekaboo dwarf allowed us to substantially improve the accuracy of its direct O/H estimate, and we obtained 12+log(O/H) = 6.990.06 dex. The new data reveal that emission lines in the east region consist of two components with a velocity difference of 65 km s-1. The fainter approaching component could be related to a fast-moving WR star thrown from a cluster or a binary system. Using the HST magnitudes and colour , we identified tentative O-type and very hot candidate WO stars, which are likely the ionising stars of the studied H ii regions.
Conclusions. With the new optical spectra, the Peekaboo galaxy is confirmed as the lowest-metallicity dwarf in the LV and as a valuable object for in-depth multi-method studies. We separate its most luminous stars for follow-up ground-based brightness monitoring and spectroscopy.
@@
The ’local’ gas-rich extremely metal-poor (XMP) dwarfs are interesting objects since they challenge modern cosmological models of galaxy formation and evolution, and they allow one to check the predictions of such models for the formation and survival of objects with extreme properties. Based on the sum of their observational properties, gas-rich XMP dwarfs appear to be similar to galaxies in the early stages of evolution.
@@
The existence of such unusual, very rare, so-called very young galaxies (VYGs) was predicted in N-body simulations (Tweed et al. 2018). The VYGs are galaxies that formed the major part of their stellar mass within the last 1 Gyr. While VYGs are expected to be very rare, their proportion in simulations strongly depends on the type of dark matter employed (i.e. cold DM versus warm DM).
@@
Apart from the limitations on the cosmological models that the statistics of XMP VYGs can imply, their study is of interest in at least in two more areas. Firstly, this is due to the expected similarity of their properties to those of young galaxies in the early Universe (very gas rich and very metal poor). Being at the distances of the Local Volume (LV) and its environs, XMP VYGs allow one to obtain much deeper insights into their properties and to tie them to the processes of galaxy formation in the early Universe.
@@
The other important issue is related to stellar physics and galaxy evolution. Namely, it concerns the basic understanding of the formation and evolution of massive stars and their feedback into the interstellar medium of their parent galaxies (e.g. Eldridge, Stanway 2022). For the cosmology of the early Universe, it is especially crucial to understand the properties and the evolution of the massive stars with metallicities tens to hundreds of times lower than the solar one Z⊙. Such massive, very low Z stars can be found in only a handful of XMP dwarf galaxies within the LV and its environs (Garcia et al. 2021). If identified, such XMP massive stars can be the primary targets of spectroscopic studies using the next generation ground-based telescopes. In the meantime, the first steps in this direction have already been presented for a nearby dIrr Leo A with Z 0.03 Z⊙ by Gull et al. (2022) and for four low-metallicity galaxies, including Leo P with Z 0.025 Z⊙ by Telford et al. (2024).
@@
The Peekaboo galaxy, the optical counterpart of the Hi source HIPASS J1131–31, got its name from the fact that it was hidden within the halo of a nearby bright star (Koribalski et al. 2018). The galaxy was recently described and studied by Karachentsev et al. (2023). As that paper reports, the ”Peekaboo dwarf is a faint (MV = –11.27 mag), the Local Volume (LV) gas-rich dIrr at the TRGB -based distance of 6.8 Mpc (distance module = 29.2 mag). Its gas metallicity, derived with the direct Te method, appeared one of the lowest of the known so far (12+log(O/H) = 6.990.16 dex)”. The cited uncertainties of O/H correspond to the range of a factor of two: [Z⊙/72–Z⊙/35]). The authors also discussed the probable ”youth” of this galaxy and highlighted the abnormally small population of aged red-giant-branch stars.
@@
In this letter, using new SALT spectroscopy, we update the estimate of Peekaboo’s metallicity and substantially improve on its accuracy. We also used archive Hubble Space Telescope (HST) images and Extragalactic Distance Database (EDD)111https://edd.ifa.hawaii.edu photometry of Peekaboo’s stars to identify the hot massive stars ionising the two H ii regions and several of the brightest supergiants. The latter can be attractive targets for HST and extremely large telescope (ELT) spectral studies, as the record-low metallicity massive evolved stars.
@@
The layout of the letter is as follows. In Sect. 2, the data and analysis are briefly described. Sect. 3 presents the main results. These results are discussed in Sect. 4 and summarised in Sect. 5. Details of the data and results are shown in the appendices. The scale we adopted is 34 pc in 1 arcsec.
@@ 2.1 SALT observations and analysis
We carried out long-slit optical spectroscopic observations of the Peekaboo galaxy using the Robert Stobie Spectrograph (RSS; Burgh et al. 2003; Kobulnicky et al. 2003) installed at the Southern African Large Telescope (SALT; Buckley, Swart & Meiring 2006; O’Donoghue et al. 2006). Different Volume Phase Holographic (VPH) gratings were used with the long slit of 1.5″ by 8′ to cover the full spectral range from 3600 Å to 7250 Å. The position angle for all observations was 87.88°, and the slit position was exactly the same as shown in Figure 6 of Karachentsev et al. (2023), where a reference star was used to put the slit in exactly the same position for different observations. A complete list of details is shown in Table 1, where all dates, seeing, covered spectral range, spectral scale, and full width at half maximum (FWHM) are shown. The final spectrum taken with grating PG0900 was used as a basic one since it covers the greatest spectral range. Spectral data obtained using other gratings were processed independently, and emission lines were measured in the same spatial areas. After that, the line fluxes, measured on the spectra obtained with other gratings, were recalculated to the flux system obtained using the PG900 grating. Namely, the total fluxes of the bright lines, measured on spectra for both gratings, were used to determine the conversion factor, and then the fluxes for the weaker lines were re-calculated using this factor. Spectrophotometric standards were observed at twilight as part of the SALT standard calibration plan.
@@
Given that SALT is equipped with an atmospheric dispersion compensator (ADC), the effect of atmospheric dispersion at an arbitrary long slit is negligible. Additionally, SALT’s design incorporates the movement of the telescope’s pupil during tracking, thereby resulting in constant alteration of the telescope’s effective area. Therefore, accurate absolute photometry and spectrophotometry are not possible. However, relative flux calibration can be used. Thus, the relative energy distribution in the spectra can be obtained with SALT data.
@@
Fortunately in our research, we had the reference star on the slit and its known magnitudes in several bands from the Legacy DR10 database Li et al. (2024). Therefore, we used this star as a local standard to derive the line fluxes and luminosity of H. (See details in Appendix, Sect. B.)
@@
The primary data reduction was done with the SALT science pipeline (Crawford et al. 2010). The long-slit reduction was done with the RSS pipeline described by Kniazev (2022). The 1D spectra extraction, emission lines measurement, and calculation of chemical abundances was done in the way described by Kniazev et al. (2008). The fully reduced 2D spectra of Peekaboo are shown in Figure 2, and the most prominent lines are labelled. Since the RSS detector is a mosaic and consists of three CCDs, the positions of two gaps between these CCDs are also shown.
@@
Errors in line intensities have been propagated in the calculations of the reddening correction, electron temperatures, and densities and have been propagated to the elemental abundance errors. The observed emission line intensities, , relative to the F(H) line and the intensities, , corrected for interstellar extinction and underlying stellar absorption are presented in Table 2. The equivalent width, (H), of the H line; the calculated absorption equivalent widths, (abs), of the Balmer lines; the extinction coefficient, (H) – which is a sum of the internal extinction in Peekaboo and foreground extinction in the Milky Way; and the subsequently recalculated value are also listed there.
@@ 2.2 Hubble Space Telescope data
Hubble Space Telescope data are well described by Karachentsev et al. (2023). Here, we used the information on positions, magnitudes, and colours of the hottest stars as the potential exciting stars of the observed H ii regions. The evolved luminous stars (supergiants) are also identified with these data, and they should be valuable in more advanced studies of the lowest metallicity massive stars using upcoming facilities, including ELTs.
In Figure 1, we show two zoomed-in HST images. In the top panel, we show an image in the F606W filter, and all the stars are marked with their numbers in Table 4. These are the same set of stars as in the paper by Karachentsev et al. (2023). In the bottom panel, we present the colour HST image of the same area of Peekaboo taken from the HST archive (Prog. SNAP 15992, PI R.B. Tully).
@@ 3.1 Structure and kinematics of the main H ii regions
Analysis of the long-slit spectra revealed the presence of two distinct H ii regions. Figure 3 shows the profile of the H line along the slit based on the spectral data obtained with the PG900 grating. The figure clearly shows two regions, which we refer to as regions ‘E’ (east) and ‘W’ (west) since the slit was oriented roughly in the east-west direction across the galaxy. These regions appear as two peaks in the H profile, although they are less pronounced in the continuum profile. Region W is fainter, with an extent of about 3 arcsec, while the brighter region E extends over about 5 arcsec. We defined approximate boundaries for these regions (as shown in the figure) in order to extract 1D spectra along the slit. The aim was to minimise spectral contamination between the regions while maximising the useful signal from the emission lines. The selected region boundaries are also shown in the Figure 2 as red horizontal lines and in the top panel of Figure 1 as red circles centred on the reference star with the corresponding radii. These boundaries were applied consistently to all the acquired spectral data. The 1D spectra obtained by summing within these limits are shown in Figures 4 and 5.
A more detailed analysis of the extracted 1D spectra uncovered an interesting peculiarity. Namely, all the observed emission lines in the spectrum of region E appear as a double component. Every line exhibits an asymmetric blue wing. To test the hypothesis that all the lines represent two velocity components, we first fit each line with a single Gaussian and then with a two-Gaussian model. The related improvement of the reduced statistic was evaluated, and we found the improvement to be very significant. For various lines, it varies by a factor of 2 to 6.5. Examples of the fitting for H and [O iii]Å lines and H and H, for gratings PG900 and PG3000 are presented in Figures 6 and 7, respectively.
We calculated the radial velocities of the two components in the E region and that of the W region, taking the weighted means on the three independent measurements with grisms PG900, PG1800, and PG3000. In turn, each of these velocities was derived from the redshifts of six to seven of the strongest emission lines. The final values of Vhel for these three subsystems are as follows: East(red): 72110 km s-1; East(blue): 6563 km s-1; and West: 6976 km s-1. For further discussion, we accepted these three Peekaboo ionised gas systems, of which East(red) and East(blue) are co-spatial within the limits of our modest spatial resolution.
@@ 3.2 New estimates of O/H and other elements
The measured and corrected line fluxes of atoms H and He and ions of O, N, Ne, S, and Ar and the derived element abundances of O, N, Ne, S and Ar are presented in Tables 2 and 3. Here, we briefly summarise the obtained results.
In Sect. 3.1, we found that there are three separate subsystems on the slit. For the East(red) subsystem, the line [O iii]4363 was detected at the level of 6.5. This allowed us to directly determine the electron temperature, Te, and derive the oxygen as well as other element abundances via the direct method. We obtained its 12+log(O/H)(Te) = 6.990.06 dex. The two non-direct methods, described in Sect. D, give for the same subsystem, the O/H values within 0.02–0.03 dex of the former. For the West and the East(blue) subsystems, the line [O iii]4363 was not detected. Their O/H were derived with the ’strong-line’ method of Izotov et al. (2019). For both objects, the derived O/H is within 0.02–0.03 dex from (O/H)(Te) of the East(red).
3.3 Hubble Space Telescope data on XMP stars in the Peekaboo dwarf
In the EDD (Anand et al. 2021) there is a table with parameters of all 56 individual stars within the Peekaboo dwarf body. We sorted the stars by colour and separated the hottest stars with –0.25 (see Sect. 4.2), which are expected to ionise the surrounding gas. We then looked for the stars that fall within (or near) the borders of the E and W H ii regions. They are the stars most likely exciting these H ii regions.
There are also several supergiants in this XMP dwarf that are very interesting since they are the nearest massive evolved stars with a record-low metallicity. They represent attractive targets for spectral study with HST and next generation ELTs. (See the top panel of Fig. 1 with the HST image of Peekaboo, where the positions of these stars are indicated.) Most of the remaining stars with of 0.0–0.5 mag and M –2.5 mag probably belong to the blue helium burning type (BHeB) with ages less than 200 Myr (McQuinn et al. 2012).